OpenStudy (anonymous):

A 25000 Kg airplane takes off from rest and reaches 5000m altitude and cruises at 900km/h in 8 minutes .Find Avg Power

OpenStudy (anonymous):

@akash123 @ghazi @ajprincess

OpenStudy (anonymous):

Pavg = W/t W = Delta K.E

OpenStudy (anonymous):

W = 25000 * 250*250 / 2

OpenStudy (anonymous):

hey do your homework yourself!!! you will not get properly educated if others will do stuff for you.

OpenStudy (ghazi):

lol

OpenStudy (anonymous):

Lol....Did anyone told this when u asked any question @ndwelsh

OpenStudy (ghazi):

\[W= \Delta K.E\] now, W= 0.5* 25000*(250)^2 and divided by 8*60 will give you power

OpenStudy (anonymous):

did you read disclaimer yourself? do you do that usually? and you know why? because plutocrats and cheaters are around us. they earn money, support families, they don't do real job. screw that. SORRY FOR THE OFFTOPIC!!!! and also, I reported this question as abuse: "students shall not ask for help doing their homework. questions should not replicate those in actual task, but should point out to according excerpts from theory. this irritates my eyes as actual professor."

OpenStudy (anonymous):

Lol..seeking help is an offence???

OpenStudy (ghazi):

@Yahoo! you must describe now what you are preparing for and how many questions are there and also explain there is no homework

OpenStudy (anonymous):

you don't get it: academic tasks are INTENDED to get done alone, without help from reference book, but rather from theory book.

OpenStudy (anonymous):

is that clear?

OpenStudy (anonymous):

but @ghazi thats wat i did but Did nt get the result..)

OpenStudy (anonymous):

ghazi: that concerns you as well, my man

OpenStudy (anonymous):

explain theory, THEN provide formula.

OpenStudy (anonymous):

you just murdered another good site, created cheating tool, so easy

OpenStudy (anonymous):

what is bad about cheating?

OpenStudy (ghazi):

@ndwelsh look sir/friend or professor i don't know what you are talking about here, this guy @Yahoo! is preparing for one of the toughest engineering exam across the globe , let him ask the way he is asking and he is not doing his homework for sure

OpenStudy (ghazi):

that exam is IIT

OpenStudy (ghazi):

jee

OpenStudy (anonymous):

again, you don't get my words. ACADEMIC questions shall be answered solely without help of reference books. And by the way, your information seems improper - I do not find anything *that* serious in this question.

OpenStudy (anonymous):

These are not my home works...)

OpenStudy (ghazi):

i am not talking about question my point of concern was if he is finding it difficult he can present his solution to be examined. let him do

OpenStudy (ghazi):

and if you don't find anything serious then probably you should help